OpenStudy (anonymous):

find the equation of the perpendicular bisector of the line segment with endpoints (-7,2)(-5,-2)

OpenStudy (anonymous):

Find the slope of the line that passes through these points. Then you can find the perpendicular bisector

OpenStudy (anonymous):

m=y2-y1/(x2-x1)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

What's the slope

OpenStudy (anonymous):

not sure, this question timed out on me, now the end points are (-5,5) (-1,-3)

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

m=(-5--1)/(5--3)

OpenStudy (anonymous):

m=-5+1/5+3=-4/8=-1/2 The slope is -1/2 A perpendicular line will have a slope that is the opposite reciprocal or 2.

OpenStudy (anonymous):

y=mx+b -3=2(-1)+b -3=-2+b -1=b Your equation perpendic...is y=2x-1

OpenStudy (anonymous):

You can follow this process for future problems

OpenStudy (anonymous):

thank you